# On the mechanics of inhaled bronchial transmission of pathogenic microdroplets generated from the upper respiratory tract, with implications for infection onset

### v1

@@ Abstract
Could the microdroplets formed by viscoelastic stretching and break-up of mucosal liquids in the upper respiratory tract (URT), when inhaled further downwind through the airway, explain the brisk pace at which deep lung infections emerge following the onset of initial infection at the URT? While it is well-established that particulates inhaled from outside can possibly penetrate to the lower airway only if they are smaller than approximately 5 $\mathrm{\mu m}$, the fate of particulates (many exceeding 5-$\mathrm{\mu m}$ in diameter) that are sheared away from the intra-URT mucosa during inhalation remains an open question. These particulates predominantly originate in the region encompassing the nasopharynx, the oropharynx, and the laryngeal chamber with the vocal folds. To resolve the posed question, this study considers a computed tomography-based three-dimensional anatomical airway reconstruction and isolates the tract from the laryngeal vocal fold region (superior to the glottis), mapping the entire tracheal cavity and concluding at the second generation of the tracheobronchial tree in the lower airway. Through the delineated geometry, airflow simulation is conducted using the Large Eddy Simulation scheme to replicate relaxed inhalation at a rate of 15 L/min. Against the ambient air flux, representative numerical experiments have been performed to monitor the transport of liquid particulates (aerosols/droplets) with diameters ranging from $\mathrm{1~{}\mu m}$ to $\mathrm{30~{}\mu m}$, bearing physical properties akin to aerosolized mucus containing embedded virions. The full-scale numerically predicted transmission trends to the lower airway were found consistent with the findings from a reduced-order mathematical model that conceptualized the impact of intra-airway vortex instabilities on local particle transport through use of point vortex idealization in an anatomy-guided two-dimensional potential flow domain. The results collectively demonstrate markedly elevated trends of deep lung penetration by the URT-derived particulates, even if they are as large as 10 $\mathrm{\mu m}$ and 15 $\mathrm{\mu m}$. The high viral load physically carried by such droplets to the bronchial spaces could mechanistically explain the accelerated seeding of infection in the lungs.
@@ I Introduction
When inhaled air sweeps past the mucociliary coating of the upper respiratory tract (URT), it is plausible to hypothesize that the interfacial interactions would lead to localized viscoelastic stretching and break-up of mucosal layers resulting in the formation and release of microdroplets, which would then be pushed downwind by the air streamlines. For instance, it is key to note the prominent instability effects when the viscoelastic layer (mucus) resting on a viscous fluid film (serous fluid) is exposed to the incoming airflow [1]. The ansatz is also supported by the established literature that has examined the reverse process: exhalation [2, 3, 4, 5]. The expiratory transport regimes, even during silent breathing, emit liquid particulates that span a wide range of length scales [6]. From such paradigm, if we now pivot our attention to inhaled (i.e., into the airway) transport, some immediate questions that come forth concern the fate of the intra-URT particulates generated during inhalation and their probable relevance in progressive disease transmission, especially to the deep lungs. On that note, the present study, through numerical experiments performed in a three-dimensional anatomical airway reconstruction and with simulation-informed reduced-order analytical modeling in an anatomy-inspired two-dimensional channel, attempts to answer the following main queries:
@@
    When inhaled from outside, it is well-known that the dominant inertial motion of the particulates larger than 5 $\mathrm{\mu m}$ ensures that hardly any of them, albeit potentially none at all, would penetrate to the lower airway. But is such conclusion also true for particulates of similar sizes generated through mucus separation from the URT surface during inhalation?
@@
    What is the viral load transmitted by particulates sheared away from the URT, that may manage to penetrate to the deep lungs, and more specifically, how does the viral transmission trend correlate with particulate sizes?
This study addresses $Q_{1}$ through full-scale numerical tracking of inhaled constituents inside an anatomically accurate airway geometry with a supporting reduced-order mathematical model of the system. Next for $Q_{2}$, it connects the fluid dynamics findings with virological parameters, such as the sputum viral concentration [7] for a representative pathogen (SARS-CoV-2), and critiques the projected viral load transmitted to the bronchial pathways in light of the verified infectious dose of the pathogen [8, 9]. The resulting translational analysis helps assess the proclivity for deep lung infection driven by inhalation of aerosolized intra-URT mucosal remnants. It is worth noting that this in silico approach, rooted in the underlying respiratory flow physics as discussed here, is agnostic to any virus specificity and is potentially extensible to other respiratory pathogens by integrating the fluid mechanics outcomes of inhaled transport with appropriate virological, epidemiological, and immunological data. In that context, this work could advance an emerging niche sub-field that brings together classical fluid dynamics and virology [10, 11, 12]. For select recent relevant studies on deep lung kinetics and pathophysiology, the reader may refer to [13, 14, 15, 16, 17].
@@
Figure 1: (a) A demonstrative cartoon of the human respiratory system, encompassing the upper respiratory tract, the mouth, and the lower respiratory tract, extending till the deep lungs. The visual is adopted with copyrights’ permission from the Getty Images®. (b-c) Sample computed tomography (CT) imaging-based reconstruction of an adult human airway. It serves as a three-dimensional anatomically realistic equivalent of the cartoon in panel (a), with regions included till generation 2 (tertiary bronchi) of the tracheobronchial tree. The domain within the red box, i.e., panel (c), is isolated for the numerical experiments on inhaled downwind transport of microdroplets generated from the upper respiratory tract. The isolated region is additionally shown in panel (d) for the sagittal view with anatomical demarcations and in panel (e) for the coronal view. The symbol $g$ signifies the gravity direction in the numerical simulations and the subsequent analytical framework, with $x$, $y$, and $z$ defining the spatial orientation of the test cavity. Box (c) additionally highlights the precise length scale of the geometry.
II Methods
II.1 Numerical modeling of inhaled transport of microdroplets generated in the upper respiratory tract
@@
As is the case for inhaled transport of microdroplets (also referred to as particulates, or equivalently, simply as particles in this exposition) through the respiratory cavity, in numerical simulations involving the dispersion of small particles under dilute conditions—the common approach assumes one-way coupling. This reduction implies that while the airflow continuum carries the particulates, the impact of such particulates on the underlying flow regimes could be disregarded. Consequently, the ambient airflow field is initially resolved, and the flow outcomes are subsequently employed to numerically solve the relevant particle equations of motion.
@@
II.1.1 Anatomical geometry reconstruction and spatial discretization
@@
From de-identified, high-resolution, medical-grade computed tomography (CT) imaging, this study has first reconstructed a complete three-dimensional adult respiratory airway [18]; see Fig. 1a-c. The extraction warranted a radio-density thresholding between $-1024$ to $-300$ Hounsfield units [19, 20] to capture the airspace from the CT slices. Aiming to address $Q_{1}$ and $Q_{2}$ (posed in §I), the reconstruction has then been digitally redacted to focus on the space mapping the vocal fold region of the larynx (space with highest concentration of liquid particulates formed from intra-URT mucosal break-up) along with the distal laryngeal chamber, the trachea, and the lower airway extending till generation 2 of the tracheobronchial tree, with primary and tertiary bronchi, followed by the respiratory bronchiolar outlets as entry to the deeper recesses of the lungs. See Fig. 1d-e for the anatomical test geometry with the described regions marked out. To prepare the domain for numerical simulations, the test cavity was spatially segregated into 2.5 million graded, unstructured, tetrahedral elements, along with four layers of pentahedral cells (with 0.025-mm height for each cell and an aspect ratio of 1.1) extruded at the airway cavity walls [21] to resolve the near-wall particulate dynamics; see Fig. 2.
Figure 2: The black box in panel (a) marks the location of the demonstrated mesh region mapping a sagittal cross-section running mid-way through the test cavity (shown henceforth in Fig. 3). Panels (b) and (c) present multi-scale visuals of the mesh. In specific, panel (c) highlights the four layers of pentahedral cells used for near-wall refinement to ensure reliable numerical tracking of particle deposition.
@@
While using the LES scheme, a finer mesh refinement typically leads to more precise outcomes. For instance, at the upper limit, achieving results akin to Direct Numerical Simulation (DNS) is possible if the grid dimensions are smaller than the Kolmogorov scale, $\mathcal{K}$, which is defined as follows [22]:
@@
with $\varepsilon$ as the turbulence dissipation rate and $\nu$ being the fluid kinematic viscosity. Another length scale of note is the Taylor scale, $\lambda$, which typically exceeds $\mathcal{K}$ and is defined as
@@
with $k$ being the turbulence kinetic energy. From the simulation data of the inhaled airflow field from this study, it is seen that both $\lambda$ and $\mathcal{K}$ collapse to $\mathcal{O}(10^{-4})$ m, while the mean grid scale is also $\rightarrow\mathcal{O}(10^{-4})$ m, suggesting that the test grid has been sufficiently resolved for precise estimation of the transport parameters.
II.1.2 Numerical simulation of inhaled airflow
We have employed the LES approach to numerically model the inhaled airflow, with eddies exceeding the grid scale explicitly resolved, whereas those that are smaller than the grid scale are approximated. Specifically, fluctuations below the grid size, referred to as subgrid scales, are filtered out, and their impact on larger scales is replicated through modeling. Assuming incompressible and isothermal conditions for the inhaled air flux, the filtered continuity and Navier–Stokes equations are respectively as follows:
@@ II.1.2 Numerical simulation of inhaled airflow
Applying the LES scheme as described above, the inhaled airflow through the anatomical airspace was replicated for 15 L/min inhalation rate [27]. See Fig. 1d for the pressure inlet face at the location of vocal folds in the larynx and for the pressure outlets at the distal ends of the reconstructed generation 2 bronchial tubes. Enforcing no slip (i.e., zero velocity) boundary condition at the airway walls, the pressure gradient-driven simulation used time-steps of 0.0002 s [25], which is smaller than the Kolmogorov time scale = $\left(\nu/\varepsilon\right)^{1/2}$ [22], for a flow solution time of 0.35 s. The simulation was executed on a segregated solver with pressure-velocity coupling and second-order upwind spatial discretization. The eventual solution convergence was monitored by minimizing the mass continuity residual to $\mathcal{O}(10^{-3})$ and the velocity component residuals to $\mathcal{O}(10^{-6})$. Also, considering the warmed-up state of inhaled air passing through the respiratory pathway, the air density $\rho$ was set at 1.204 kg/m3 in the simulations, with 18.218$\times$10-6 m2/s as its kinematic viscosity $\nu$.
II.1.3 Numerical experiments for particulate transport
Against the solved airflow field, the particulates assumed to be formed from intra-URT mucus break-up were tracked from the vocal fold region of the larynx (see the marked inlet face in Fig. 1d). The air-particle phases were one-way coupled with the particles being impacted by the ambient flow field; the underlying flow domain was considered quasi-steady while evaluating the particle transport parameters. Lagrangian-based inert discrete phase model, with a Runge-Kutta solver, was used to numerically integrate the particle transport equation:

### v2

@@ Abstract
Could the microdroplets formed by viscoelastic fragmentation of mucosal liquids within the upper respiratory tract (URT) explain the brisk onset of deep lung infection following initial URT infections? Generally, particulates, inhaled through the nostrils and therefore navigating the intricate topography of the anterior nasal cavity, can efficiently reach the lower airway only if they are small enough, typically $\lesssim 5\,\mathrm{\mu m}$. However, the fate of larger particulates, many exceeding 5-$\mathrm{\mu m}$ in diameter, that are sheared from the initial infection sites along the intra-URT mucosa during inhalation remains unresolved. These particulates originate primarily from the nasopharynx, oropharynx, and the laryngeal chamber containing the vocal folds. To investigate, this study employs a computed tomography-based three-dimensional anatomical airway reconstruction, isolating the tract from the larynx and mapping the tracheal cavity through to the third generation of the tracheobronchial tree; constituent transport across the distal bronchial outlets is also recorded to assess deep lung penetration. Within the defined geometry, airflow simulations are conducted with the Large Eddy Simulation scheme to replicate relaxed inhalation at 15 L/min flow rate. Against the ambient air flux, numerical experiments are performed to monitor the transport of particulates (aerosols/droplets) with diameters $\mathrm{1-30~{}\mu m}$, bearing physical properties akin to aerosolized mucus with embedded virions. The full-scale numerical transmission trends are consistent with findings from our reduced-order mathematical model that conceptualizes the influence of intra-airway vortex instabilities on local particle transport through point vortex idealization in an anatomy-guided two-dimensional potential flow domain. The results collectively demonstrate markedly elevated lower airway penetration by URT-derived particulates, even by those as large as 10 and 15 $\mathrm{\mu m}$. The high viral load, often exceeding the pathogen-specific infectious dose, carried by such droplets into the bronchial spaces of the sample airway, provides a plausible mechanistic explanation for the accelerated seeding of secondary lung infection.
@@ I Introduction
When inhaled air sweeps past the mucociliary coating of the upper respiratory tract (URT), interfacial interactions lead to localized viscoelastic stretching and breakup of mucosal layers resulting in the formation and release of microdroplets [1, 2, 3], which could then be pushed downwind by the airflow streamlines. Prominent instability effects when the viscoelastic layer (mucus) resting on a viscous fluid film (serous fluid) is exposed to the incoming airflow are to be noted in this context; e.g., see [4]. The mechanism is analogous to numerous observations from the reverse process: exhalation [5, 6, 7, 8]. The expiratory transport regimes, for intense respiratory events [9] and even during silent breathing [10], often involve mucus fragmentation and subsequent emission of liquid particulates spanning a wide range of length scales. From that paradigm, if we pivot our attention to inhaled (i.e., into the airway) transport, some immediate questions that come forth concern the fate of the intra-URT particulates generated during inhalation and their probable relevance in progressive disease transmission, especially to the deep lungs. On that note, the present study, through numerical experiments performed in a three-dimensional anatomical airway reconstruction and with simulation-informed reduced-order analytical validation in an anatomy-inspired two-dimensional channel, attempts to answer the following main queries:
@@
    If inhaled from outside and consequently navigating the complex anterior nasal topography, the dominant inertial motion as well as gravitational impaction for particulates larger than 5 $\mathrm{\mu m}$ may, in general, prevent their penetration to the lower airway [11, 12]. However, is this also true for particulates of similar sizes shed through mucus separation at the back of the nasal passage from the URT sub-sites like the nasopharynx, oropharynx, and from around the vocal folds during inhalation? Hypothetically, with such particulates still being airborne at the larynx (see Fig 1), the relatively straight structural shape of the downwind tract up to the tracheal base may facilitate their penetration into the lower airway.
@@
    What is the viral load transmitted to the lower airway by particulates originating from the infected intra-URT tissue surfaces? Could this mechanism, in general, explain the rapid onset of lung infection following the initial infection and the emergence of symptoms at URT sub-sites, such as the nasopharynx? First proposed (to the author’s knowledge) in [13], the veracity of such a mechanism can dispel the caveat concerning time-scale inconsistencies whereby attributing the rapid downwind progression of infection solely to tissue-level replication of the invading pathogen could be a stretch.
This study addresses $Q_{1}$ through full-scale numerical tracking of inhaled constituents inside a representative anatomically accurate airway geometry with a supporting reduced-order anatomy-guided mathematical model of the system. The anatomical test geometry extends up to the third generation of bronchial branching, with cavity outlets leading into finer respiratory bronchioles; see Fig 1. To generate a summative assessment of bronchial deposition alongside deep lung penetration of URT-derived particulates, this study extracts and combines: (a) the simulated bronchial deposition percentages up to generation 3; and (b) the percentage of particulates progressing further downwind through the distal bronchial outlets into the deeper respiratory bronchioles. To note, the term deep lungs refers to the lower regions of the respiratory system where gas exchange primarily occurs. This domain includes the bronchioles, alveolar ducts, and alveoli [14]. A detailed analysis of, say, the alveolar deposition trends in the deeper regions of the lungs lies beyond the scope of the present methodology. Using physiological modeling in a test respiratory domain, this paper exclusively assesses the lower airway penetration potential for URT-derived pathogenic microdroplets and how the mechanism could be correlated to secondary infection onset. However, for supplementary reading, comprehensive bronchial deposition profiles for (mostly pharmaceutical) particles administered nasally and/or orally from outside can be found in several excellent studies; e.g., see [15, 16, 17, 18, 19, 20, 21].
@@
Figure 1: Defining the physiological domain. (a) A demonstrative cartoon of the human respiratory system, encompassing the upper respiratory tract, the mouth, and the lower respiratory tract, extending till the deep lungs. The visual is adopted with a perpetual license agreement from the Getty Images®. (b) Sample computed tomography (CT) imaging-based reconstruction of an adult human airway. It serves as a three-dimensional anatomically realistic equivalent of the cartoon in panel (a), with regions included till generation 3 (tertiary bronchi) of the tracheobronchial tree. For confirmation, note that there are 8 distal outlets (see panels (b)-(d)), implying $2^{G_{n}}$ = 8, where $G_{n}=3$ is the generation number, considering two-way bifurcation for each bronchial tube at every transition [22]. The domain within the red box is isolated for the numerical experiments on inhaled downwind transport of microdroplets generated from the upper respiratory tract. The isolated region is additionally shown in panel (c) for the sagittal view with anatomical demarcations and in panel (d) for the coronal view. The symbol $g$ signifies the gravity direction in the numerical simulations and the subsequent analytical framework, with $x$, $y$, and $z$ defining the spatial orientation of the test cavity. Panel (b) additionally highlights the geometric length scale.
Next for $Q_{2}$, the fluid dynamics findings on particulate penetration are integrated with virological parameters, such as the sputum viral concentration [23] for a representative pathogen (SARS-CoV-2), and the projected viral load transmitted to the bronchial pathways is compared against the verified infectious dose of the pathogen [24, 25]. The resulting translational analysis helps assess the proclivity for lower airway infection that is driven by inhalation of aerosolized intra-URT mucosal remnants. It is worth noting that this in silico approach, rooted in the underlying respiratory flow physics as discussed here, is agnostic to any virus specificity and potentially extensible to other respiratory pathogens by integrating the fluid mechanics outcomes of inhaled transport with appropriate virological and immunological data. In that context, this work can contribute toward advancing an emerging niche sub-field that brings together classical fluid dynamics and virology [26, 27, 28, 29, 25, 30]. For select recent relevant studies on deep lung kinetics and pathophysiology, the reader may refer to [31, 32, 33, 13, 34]; additionally, one may peruse [35, 36] for engaging accounts of airway flow physics, its potential impact on lung function, and the pertinent open problems.
II Materials and methods
Numerical modeling of inhaled downwind transport of URT-derived microdroplets
@@ II Materials and methods
As is the case for inhaled transport of microdroplets (also referred to as particulates, or equivalently simply at times as particles in this exposition) through the respiratory cavity, in numerical simulations involving the dispersion of small particles under dilute conditions—the common approach assumes one-way coupling. This reduction implies that while the airflow continuum carries the particulates, the impact of such particulates on the underlying flow regimes could be disregarded. Consequently, the ambient airflow field is initially resolved, and the flow outcomes are subsequently employed to numerically solve the relevant particle equations of motion.
@@
II.0.1 Anatomical geometry reconstruction, spatial discretization, and mesh sensitivity analysis
@@
From de-identified, high-resolution, medical-grade computed tomography (CT) imaging, this study has first reconstructed a complete three-dimensional adult respiratory airway [37]; see Fig 1a-b. The extraction warranted a radio-density thresholding between $-1024$ to $-300$ Hounsfield units [38, 39] to capture the airspace from the CT slices. Aiming to address $Q_{1}$ and $Q_{2}$ (pitched in the introduction), the reconstruction has then been digitally redacted to focus on the space mapping the vocal fold region of the larynx (space with highest concentration of liquid particulates formed from intra-URT mucosal fragmentation) along with the distal laryngeal chamber, the trachea, and the lower airway extending till generation 3 of the tracheobronchial tree, incorporating the primary, secondary, and tertiary bronchi, followed by the respiratory bronchiolar outlets as entry to the deeper recesses of the lungs. See Fig 1c-d for the redacted test geometry with the described regions marked out. To prepare the domain for numerical simulations, a grid refinement analysis was conducted with the test cavity being spatially segregated into 0.5, 1.0, 1.5, 2.0, 2.5, and 3.0 million graded, unstructured, tetrahedral elements, along with four layers of pentahedral cells (with 0.025-mm height for each cell and an aspect ratio of 1.1) extruded at the airway cavity walls [40] to resolve the near-wall particulate dynamics; see Fig 2a-h. The variables assessed in the sensitivity study included the resistance $\mathcal{R}$ (in Pa.min/L) to the simulated inhaled airflow, calculated as $|\Delta P|/Q$, where $|\Delta P|$ in Pa represents the inlet-to-outlet pressure gradient driving the flow and $Q$ denotes the volumetric flux in L/min; the area-weighted average airflow velocity magnitude $V_{o}$ (in m/s) at the outlet surfaces of the geometry; and the transmission efficiency $\eta$ (in %) within the bronchial pathways for representative particulate sizes of 1, 5, 10, and 20 $\mathrm{\mu m}$. The deposition and penetration efficiency $\eta$ quantifies the cumulative bronchial transmission trend for each particulate size—summing the deposition rates at the primary, secondary, and tertiary bronchi, together with the penetration rate through the distal bronchiolar outlets into the deeper lung regions (see Fig 1c-d). The fluctuation trends of the tracked variables are shown in Fig 2i-k. Specifically, the simulation results yield:
$\sigma(\mathcal{R})_{\rm{All}}$ $=0.0051~{}\rm{Pa.min/L},$ $\sigma(\mathcal{R})_{\rm{3}}$ $=0.0041~{}\rm{Pa.min/L},$ (1)
$\sigma(V_{o})_{\rm{All}}$ $=0.0085~{}\rm{m/s},$ $\sigma(V_{o})_{\rm{3}}$ $=0.0026~{}\rm{m/s},$
$\sigma(\eta)_{\rm{All}}$ $=7.94\%,$ $\sigma(\eta)_{\rm{3}}$ $=0.44\%,$
where $\sigma(\cdot)_{\rm{All}}$ denotes the standard deviation across all six grids and $\sigma(\cdot)_{\rm{3}}$ denotes the standard deviation of the simulated data from the last 3 grids (i.e., cases with 2.0, 2.5, and 3.0 million unstructured tetrahedral cells). Based on the asymptotic convergence trend observed in these final three cases, the intermediate 2.5-million cell grid resolution was selected for the overall study. The choice is consistent with detailed studies on grid convergence and computational stability for simulating physiologically realistic airflow and particle deposition in the human respiratory system; e.g., see [41].
Figure 2: Assessing the numerical sensitivity. Panel (a) marks the location of the demonstrated unstructured tetrahedral meshes spatially refined with the following cell counts (in million): (b) 0.5, (c) 1.0, (d) 1.5, (e) 2.0, (f) 2.5, (g) 3.0. Panel (h) highlights the four layers of pentahedral cells used for near-wall refinement. Panels (i)-(k) present the computed flow and particulate transport variables across the six mesh resolutions: (i) resistance $\mathcal{R}$ to inhaled airflow, (j) area-weighted average airflow velocity $V_{o}$ at the outlet faces, and (k) deposition (or, penetration) rate $\eta$ (%) in bronchial pathways for select particle sizes of 1, 5, 10, and 20 $\mu m$. The circle sizes to the right of panel (k) represent the number of tracked particulates in each simulation, scaled proportionally: red = 597, orange = 831, ocher = 1208, green = 1459, dark green = 1622, and blue = 1954 tracked particulates per size. Tracked particle counts equal the number of surface elements at inlet, directly proportional to the spatial resolution of the respective mesh.
@@
Notably, while using the LES scheme, a finer mesh refinement typically leads to more precise outcomes. For instance, at the upper limit, achieving results akin to Direct Numerical Simulation (DNS) is possible if the grid dimensions are smaller than the Kolmogorov scale, $\mathcal{K}$, defined as [42]:
@@
with $\varepsilon$ as the turbulence dissipation rate and $\nu$ being the fluid kinematic viscosity. Another length scale of note is the Taylor scale, $\lambda$, which typically exceeds $\mathcal{K}$ and is defined as:
@@
with $k$ being the turbulence kinetic energy. From the simulation data of the inhaled airflow field, it is seen that both $\lambda$ and $\mathcal{K}$ collapse to $\mathcal{O}(10^{-4})$ m, while the mean grid scale is also $\rightarrow\mathcal{O}(10^{-4})$ m, suggesting that the test grid has been sufficiently resolved for reliable estimation of the transport parameters.
II.0.2 Numerical simulation of inhaled airflow
We have implemented the LES approach to numerically model the inhaled airflow, with eddies exceeding the grid scale explicitly resolved, whereas those that are smaller than the grid scale are approximated. Specifically, fluctuations below the grid size, referred to as subgrid scales, are filtered out, and their impact on larger scales is replicated through modeling. Assuming incompressible, isothermal conditions for the inhaled air flux, the filtered continuity and Navier–Stokes equations are respectively:
@@ II.0.2 Numerical simulation of inhaled airflow
Applying the LES scheme as described above, the inhaled airflow through the anatomical airspace was replicated for 15 L/min inhalation rate, conforming with normal relaxed breathing conditions [47]. See Fig 1c for the pressure inlet face at the location of vocal folds in the larynx and for the pressure outlets at the distal ends of the reconstructed generation 3 bronchial tubes. Enforcing no slip (i.e., zero velocity) boundary condition at the airway walls, the pressure gradient-driven simulation used time-steps of 0.0002 s. The latter was chosen based on reported findings [45] on the time-step size warranted to fully resolve the unsteady turbulent airflow field in a realistic upper airway model, it being smaller than the Kolmogorov time scale = $\left(\nu/\varepsilon\right)^{1/2}$ [42], for a flow solution time of 0.35 s. The simulation was executed on a segregated solver with pressure-velocity coupling and second-order upwind spatial discretization. For the transient formulation, a bounded second-order implicit scheme was employed, to strike a balance between accuracy (due to the second-order formulation), stability (from the implicit approach), and boundedness (to prevent non-physical oscillations). The eventual solution convergence was monitored by minimizing the mass continuity residual to $\mathcal{O}(10^{-3})$ and the velocity component residuals to $\mathcal{O}(10^{-6})$. Also, considering the warmed-up state of inhaled air passing through the respiratory pathway, the air density $\rho$ was set at 1.204 kg/m3 in the simulations, with 15.16$\times$10-6 m2/s as its kinematic viscosity $\nu$.
II.0.3 Numerical experiments for particulate transport
Against the solved airflow field, the particulates formed intra-URT were tracked from the vocal fold region of the larynx (see the marked inlet face in Fig 1c). The air-particle phases were one-way coupled with the particles (assumed spherical) being impacted by the ambient flow field; the underlying flow domain was considered quasi-steady while evaluating the particle transport parameters. Lagrangian-based inert discrete phase model, with a Runge-Kutta solver, was used to numerically integrate the particle transport equation: